# The expected value of sample information calculations for external validation of risk prediction models

### v2

@@ Results
We used data from GUSTO-I, a clinical trial of multiple thrombolytic strategies for acute myocardial infarction (AMI), as our case study26. This dataset has been widely used for methodological research in predictive analytics27. All analyses are conducted in R28.
For the present case study, we use a similar set-up as our previous work, with a risk prediction model for 30-day mortality developed using the non-US sample of this dataset, and validated in the US sub-sample. The risk prediction model is a logit model fitted to the entire non-US subset (sample size 17796), with the resulting equation:
@@
Similar to our previous work, we initially assume we have access to only $n=500$ individuals from the US sub-sample as the source of current information. We randomly selected 500 individuals, without replacement, from the entire 23034 observations in the US sub-sample of GUSTO-I. Later, we will use more observations from the US sub-sample to study how EVSI behaves with the amount of current information. The prevalence of the outcome in the development sample (entire non-US sub-sample) was 0.0723; in the current validation sample it was 0.0860, while in the entire US sub-sample it was 0.0679. We consider the range of relevant thresholds to be 1%-10%.
We start by drawing the decision curve at this range of thresholds, as shown in Figure 1. Panel (a) shows the decision curve, and panel (b) shows the difference between the NB of the model and the best alternative default strategy ($NB_{1}-\max(NB_{0},NB_{2})$) along with its 95% confidence interval (CI, based on the percentile methods using 10,000 bootstraps). The model had higher expected NB than the alternative strategies at the entire range of thresholds of interest.
(a) Decision curve
(b) Incremental NB of the model v. best alternative strategy
Figure 1: Decision curve (a) and incremental net benefit of the model v. the best alternative strategy (b). Dashed lines are 95% confidence intervals
Next, we draw the EVPI curve. EVPI and EVSI computations are based on 10^6 Monte Carlo simulations using the aggregate beta-binomial model, with a flat $\text{Beta}(1,1)$ prior on each component of $\theta$. The EVPI at 0.01 threshold was 0.00039. At the 0.02 threshold it was 0.00125. The EVPI was 0 at 0.05 and 0.10 thresholds (so EVSI for any $n^{*}$ will also be 0). We therefore focus on VoI calculations at 0.01 and 0.02 thresholds.
The EVPI curve is presented in Figure 2. EVPI quantifies the expected NB loss per decision (every time the model might be used). The overall NB loss due to uncertainty is therefore affected by the expected number of times the medical decision will be made29. In scaling VoI quantities to the population, we apply a similar approach as in our previous work11: a decision-making body in charge of making a national recommendation can consider the model to be applicable to all AMI cases in the US (the target population of the validation sample). There are around 800,000 AMIs in the US every year30, and a guideline development panel in charge of recommending risk stratification for AMI management can consider all those instances relevant. The scaled VoI values are provided in true positive units on the second Y-axis of the EVPI figure. At the 0.02 threshold, a future validation study can have a maximum benefit equal to gaining 1001 net true positive (individuals who will die within 30 days and are correctly identified as high risk) per year, or avoiding 49069 extra false positive cases per year (individuals who will not die within 30 days.
@@
The EVSI curves are shown in Figure 3. Generally, it is expected that EVSI will increase with a larger future study, and will asymptote to EVPI, a pattern that is obvious for both thresholds. The ‘diminishing return’ pattern is also obvious for both thresholds: the expected NB gain steeply rises at small samples and plateaus as $n^{*}$ grows. Similar to EVPI, population scaling is applied to EVSIs and is presented as the second Y-axis of the EVSI curve. At 0.02 threshold, a future study of size $n^{*}=500$ has a per-decision EVSI value of 0.00100, corresponding to population value of 804 in true positive cases gained, or 39381 in false positive cases averted.
(a) $z=0.01$
(b) $z=0.02$
Figure 3: EVSI for the case study
Figure 4 provides the results of brief simulation studies that demonstrate how EVSI changes as a function of the amount of current information, represented by the sample size of $\mathbf{d}$, the current validation data from which $P(\theta|\mathbf{d})$ is constructed. Starting from a value of $500$, we doubled the sample size of the current validation sample, specified the distributions for $\theta_{p}$, $\theta_{se}$, and $\theta_{sp}$ based on this sample, and performed EVSI calculations as above, until reaching $n=8,000$. Results are provided for both 0.01 and 0.02 thresholds and are the average of 100 independent simulations, in each a new sample (with replacement) from the US subset of GUSTO-I was obtained as the current validation data (100 simulations were considered adequate based on inspecting the standard error of the Monte Carlo simulations).
As expected, the higher the amount of current information (the larger the $n$), the lower the gain in NB by conducting future studies. When the current validation sample is small, the initial gain in NB by conducting future research can be very large, as evidenced by the steep slope of the EVSI curve. This is not the case when the current information is relatively robust (e.g., based on $n$=8,000, which includes on average 544 events). The comparison between the two thresholds is also informative, as it demonstrates that the amount of gain from future sample is highly dependent on the threshold, even in relative terms. For example, at $z=0.01$, the EVSI curve when current information is based on $n$=1,000 (dotted blue line) is only slightly lower than the EVSI curve when $n$=500 (dotted black line). At $z=0.02$, EVSI when current information is based on $n$=1,000 (solid blue line) is significantly lower than when the current information is based on $n$=500 (solid black line), across the entire range of future sample sizes.
Figure 4: The expected value of sample information (EVSI) as a function of the sample sample size of the future study ($\mathcal{D^{*}}$) for various levels of current information, represented by the sample size of the current study ($n$) from which $P(\theta)$ is constructed
@@ Discussion
How can EVSI analysis inform study design in predictive analytics? Conducting a validation study is an investment in resources that will generate further information on the performance of a clinical prediction model in a target population. The decision whether to undertake a validation study should ultimately hinge on whether the information gained from such a study is worth the required investments. The EVSI, when scaled to the population, determines the expected return on investment in NB unit. Ultimately, such return should be contrasted against the efforts and resources required for such a study. In decision-analytic (health policy) modeling, EVSI is typically in net monetary units, and when scaled to population, can be compared with the budget of a planned data collection activity. The optimal sample size will be one that maximizes the difference between population EVSI and study costs29. The NB for risk prediction models, on the other hand, is in net true or false positive units, and as such cannot directly be compared with the budget of a validation study. One can always embark on full decision analysis to translate all outcomes to net monetary units, but this will likely require sophisticated decision modeling and context-specific assumptions on long-term outcomes, a process that might take significant amount of time and require further data collection (e.g., to obtain utility weights for outcomes). To us, a main reason for the vast popularity of decision curve analysis is that it provides an assumption-free, reproducible method for NB calculation based on the very same data that are used for studying model calibration and discrimination. We think VoI analysis during model development and validation should generally keep the same spirit. A full decision analysis should be relegated to after an impact analysis has measured the resource-use implications of implementing the model. This, however, means the decision rules for determining the optimal sample size of development and validation studies based on the VoI framework would be different than those used in health policy analysis. This paper deliberately stayed away from proposing such rules, and instead focused on defining concepts and proposing computation methods for EVSI. Proposing such decision rules is detached from EVSI calculations and deserves its own airing.
@@
There are multiple areas of further inquiry. The EVSI framework should also be applied to the development phase of prediction models. This can guide the investigator on whether further development, or moving to validation, should be prioritized. We mainly focused on NB loss due to sampling uncertainty. However, there are several sources of uncertainty, such as whether our existing information on the performance of the model is directly applicable to the target population, or if predictors and outcome are measured with the same quality between the study and usual practice. The comparative statistical and computational performance of the EVSI computations algorithms, and the adequacy of a given number of simulations for each algorithm, should be evaluated in dedicated studies. We also do not claim the algorithms we proposed are the only ones that can be used for EVSI computations. Other algorithms, such as those based on central limit theorem31,32, can prove useful. VoI analysis in decision modeling has received significant boost in computational speed in recent years due to the arrival of algorithms based on non-parametric regression modeling33,34. This approach can facilitate VoI analysis in risk prediction as well. The EVSI defined in this work is for a single, homogeneous target population, and does not consider heterogeneous settings. Consider the hierarchical Bayesian tri-variate model by Wynants et al for NB estimation in multi-center validation studies19 and imagine a new center is interested in validating this model in their local setting. The posterior predictive distribution from the hierarchical model for a new center (potentially combined with any local information on the performance of the model) can act as ‘current information’, enabling EVSI calculations. This will be valid and informative insofar as the expected NB gain in that center is concerned. However, in a hierarchical Bayesian framework, information learned from one cluster propagates to others, and thus conducting a validation study in a new center results in change in the distribution of NB in other centers. Such NB gain from learning across clusters in hierarchical settings is not reflected in the EVSI computed for this particular setting. VoI metrics and corresponding computation algorithms for multi-center studies should be developed separately. Further, during external validation, often a secondary aim is to update the model if its performance turns out to be sub-optimal35. Such model revision can take different levels of complexity (from intercept correction to re-estimating regression coefficients36) and one might be interested in the expected yield of a given sample size for such model revision. This will get connected to the VoI concepts for model development. As stated earlier, how VoI metrics for prediction models should inform objective functions for determining the optimal sample size should be debated by the community in the hope of generating consensus and best practice standards.
@@
There is an ongoing debate on the appropriateness of conventional, frequentist metrics of uncertainty when reporting NB of a clinical prediction model3–5. The same concerns can logically be extended to frequentist method for sample size and power calculations around NB37–39. VoI methodology provides a rigorous, utilitarian response to such controversies. The toolbox of VoI methods for clinical prediction models is growing, and perhaps it is time to formalize the role of VoI in uncertainty quantification and design of empirical studies in predictive analytics.

### v3

@@ Results
The prevalence of the outcome in the development sample (entire non-US sub-sample) was 0.0723; in the current validation sample it was 0.0860, while in the entire US sub-sample it was 0.0679. The risk prediction model was of the form:
@@
The decision curve depicting the NB of the model alongside alternative strategies is presented in Figure 1. Panel (a) shows the NB, and panel (b) shows the difference between the NB of the model and the best alternative default strategy ($NB_{1}-\max(NB_{0},NB_{2})$) along with its 95% CI (based on the percentile methods using 10,000 bootstraps). The model had higher expected NB than the alternative strategies at the entire range of thresholds of interest.
(a) Decision curve. Solid blue: use the model; dashed gray: treat all; dotted gray: treat none
(b) Incremental NB of the model versus best alternative strategy
Figure 1: Decision curve (a) and incremental net benefit of the model versus the best alternative strategy (b). Dashed lines are 95% confidence intervals
Among 500 observations in the current sample, 43 experienced the outcome. Given the prior $\text{Beta}(1,1)$, this will result in $\theta_{p}\sim\text{Beta}(44,458)$. The distribution of sensitivity and specificity depends on the threshold. Taking $z=0.02$ as an example, given that 41 of the 43 individuals who experienced the outcome had predicted risk $\geq z$, and 147 of the 457 who did not experience the outcome had predicted risk $<z$, with a $\text{Beta}(1,1)$ prior, we have $\theta_{se}\sim\text{Beta}(42,3)$ and $\theta_{se}\sim\text{Beta}(148,311)$.
The EVPI curve is presented in Figure 2. The EVPI at 0.01 threshold was 0.00037. At the 0.02 threshold it was 0.00125. The EVPI was 0 at 0.05 and 0.10 thresholds (so EVSI for any $\mathit{N}$ will also be 0). We therefore focus on VoI calculations at 0.01 and 0.02 thresholds. EVPI quantifies the expected NB loss per decision (every time the model might be used). The overall NB loss due to uncertainty is therefore affected by the expected number of times the medical decision will be made36. In scaling VoI quantities to the population, we apply a similar approach as in our previous work15: There are around 800,000 AMIs in the US every year37, and a national guideline development panel in charge of recommending risk stratification for AMI management can consider all those instances relevant. The scaled VoI values are provided in true positive units on the second Y-axis of the EVPI figure. At the 0.02 threshold, a future validation study can have a maximum benefit equal to gaining 1,004 net true positive (individuals who will die within 30 days and are correctly identified as high risk) per year, or avoiding 49,188 extra false positive cases per year (individuals who will not die within 30 days but are identified as high risk).
@@
The EVSI curves are shown in Figure 3. Generally, it is expected that EVSI will increase with a larger future study, and will asymptote to EVPI, a pattern that is obvious for both thresholds. The ‘diminishing return’ pattern is also clear: the expected NB gain steeply rises at small samples and plateaus as $\mathit{N}$ grows, leaving little room for gaining NB beyond $\mathit{N}$=4,000. Similar to EVPI, population scaling is applied to EVSIs and is presented as the second Y-axis of the EVSI curve. At 0.02 threshold, a future study of size $\mathit{N}=1,000$ has a per-decision EVSI value of 0.00101, corresponding to population value of 806 in true positive cases gained, or 39,500 in false positive cases averted.
Figure 3: The expected value of sample information (EVSI) for the case study. Red: EVSI; blue (horizontal): the expected value of perfect information (EVPI). Dashed lines: $z$ (risk threshold)=0.01; solid lines: $z$=0.02
@@ Discussion
We proposed three algorithms for EVSI computations (with implementation in R as part of the evsiexval package: https://github.com/resplab/evsiexval). The bootstrap-based algorithm is applicable when previous individual-level data (e.g., from a pilot validation study) are at hand. The main advantage of this algorithm is in its flexibility, for example in mimicking expected patterns of missingness in data. Among the three algorithms, this algorithm is the one that can most readily be extended to survival outcomes (similar to the extension of the bootstrap method for inference around NB16). However, this algorithm is applicable only when individual-level data are available and can be slow. The beta-binomial algorithm is applicable when current evidence on model performance can be expressed as independent beta distributions for prevalence, sensitivity, and specificity, and is the fastest of the three algorithms. We showed that for binary outcomes and in the absence of censoring this algorithm is equal to the Bayesian bootstrap (with the added flexibility that prior information can be incorporated). The general algorithm is the most versatile one as it works with any joint distribution of $\theta$ as long as one can obtain random draws from it. As such, this algorithm can be useful when $P(\theta|\mathbf{d})$ does not have a tractable form. However, the finite size of the sample adds another layer of uncertainty, potentially requiring significantly more computation time to achieve the same numerical accuracy of the other two algorithms.
@@
How can EVSI analysis inform study design in predictive analytics? Conducting a validation study is an investment in resources that will generate further information on the performance of a clinical prediction model in a target population. The decision whether to undertake a validation study should ultimately hinge on whether the information gained from such a study is worth the required investments. The EVSI, when scaled to the population, determines the expected return on investment in NB unit. Ultimately, such return should be contrasted against the efforts and resources required for such a study. In decision-analytic (health policy) modeling, EVSI is typically in net monetary units, and when scaled to population, can be compared with the budget of a planned data collection activity. The optimal sample size will be one that maximizes the difference between population EVSI and study costs36. The NB for risk prediction models, on the other hand, is in net true or false positive units, and as such cannot directly be compared with the budget of a validation study. One can always embark on full decision analysis to translate all outcomes to net monetary units, but this will likely require sophisticated decision modeling and context-specific assumptions on long-term outcomes, a process that might take significant amount of time and require further data collection (e.g., to obtain utility weights for outcomes). To us, a main reason for the vast popularity of decision curve analysis is that it provides an assumption-free, reproducible method for NB calculation based on the very same data that are used for studying model calibration and discrimination. We think VoI analysis during model development and validation should generally keep the same spirit. A full decision analysis should be relegated to after an impact analysis has measured the resource-use implications of implementing the model. This, however, means the decision rules for determining the optimal sample size of development and validation studies based on the VoI framework would be different than those used in health policy analysis. This paper deliberately stayed away from proposing such rules, and instead focused on defining concepts and proposing computation methods for EVSI. Proposing such decision rules is detached from EVSI calculations and deserves its own airing.
@@
There are multiple areas of further inquiry. The EVSI framework should also be applied to the development phase of prediction models. This can guide the investigator on whether further development, or moving to validation, should be prioritized. We mainly focused on NB loss due to sampling uncertainty. However, there are several sources of uncertainty, such as whether our existing information on the performance of the model is directly applicable to the target population, or if predictors and outcome are measured with the same quality between the study and usual practice. The comparative statistical and computational performance of the EVSI computations algorithms, and the adequacy of a given number of simulations for each algorithm, should be evaluated in dedicated studies. We also do not claim the algorithms we proposed are the only ones that can be used for EVSI computations. Other algorithms, such as those based on central limit theorem38,39, can prove useful. VoI analysis in decision modeling has received significant boost in computational speed in recent years due to the arrival of algorithms based on non-parametric regression modeling40,41. This approach can facilitate VoI analysis in risk prediction as well. The EVSI defined in this work is for a single, homogeneous target population, and does not consider heterogeneous settings. VoI metrics and corresponding computation algorithms for multi-center studies should be developed separately. Further, during external validation, often a secondary aim is to update the model if its performance turns out to be sub-optimal42. Such model revision can take different levels of complexity43 and one might be interested in the expected yield of a given sample size for such model revision. This will get connected to the VoI concepts for model development. As stated earlier, how VoI metrics for prediction models should inform objective functions for determining the optimal sample size should be debated by the community in the hope of generating consensus and best practice standards.
@@
There is an ongoing debate on the appropriateness of conventional metrics of uncertainty when reporting NB of a clinical prediction model7–9. The same concerns can logically be extended to frequentist method for sample size and power calculations around NB5,6. VoI methodology provides a rigorous, utilitarian response to such controversies. The toolbox of VoI methods for clinical prediction models is growing, and perhaps it is time to formalize the role of VoI in uncertainty quantification and design of empirical studies in predictive analytics.